# Colored Noise in PPO: Improved Exploration and Performance Through Correlated Action Sampling

### 18 Dec 2023 v1

@@ 1 Introduction
In this paper, we extend previous work on correlated action-noise in off-policy RL to on-policy RL methods: we introduce and empirically evaluate a modification to PPO incorporating temporally correlated colored noise into the stochastic policy’s distribution. Utilizing the re-parameterization trick, we maintain a Gaussian distributed behavior while injecting correlated noise instead of uncorrelated noise. This introduces a new hyper-parameter, the noise color $\beta$ parameterizing the correlated noise.
@@ 1.1 Related Work
In off-policy methods, exploration is often achieved by perturbing the action selection process, for example by perturbing the parameters of the policy (Plappert et al. 2018) or by additively perturbing the action, i.e., by adding action noise. This action perturbation can be done, for example, using temporally uncorrelated Gaussian noise, or temporally correlated noise generated by an Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein 1930). Previous research has shown that increasing the temporal correlation of actions, i.e., switching from Gaussian noise to Ornstein-Uhlenbeck noise, tends to increase state space coverage, but an increase in state space coverage can be beneficial or harmful for policy learning; this depends on the task environment and its dynamics (Hollenstein et al. 2021, 2022). Another method for temporally correlated perturbations was introduced by Raffin, Kober, and Stulp (2021) where the “action noise” is generated deterministically by a function only dependent on the state, but the function parameters are randomly changed after a number of steps. This approach is extended by Chiappa et al. (2023) to exploit the hidden layers of the policy, to induce correlation between the action dimensions.
@@ 3.1 Evaluation Details
Environment $\beta^{*}$
Noise Color
$\beta\newline 0.5$
$p_{0.5}$
$\beta\newline 0.0$
$p_{0.0}$
$\beta\newline 1.0$
$p_{1.0}$
4 Rooms Swimmer 0.00
White
0.19
0.01
Cartpole Balance 0.00
White
0.04
0.01
Cheetah Run 0.00
White
0.14
0.01
Reacher Hard 0.00
White
0.01
0.01
UMaze Ant 0.00
White
0.24
0.01
UMaze Swimmer 0.00
White
0.16
0.01
4 Rooms Ant 0.20
0.21
0.98
0.01
Door 0.20
0.06
0.56
0.25
Cartpole Swingup 0.50
0.03
0.01
4 Rooms Point 0.75
0.97
0.03
0.51
Ball in Cup (Catch) 0.75
0.90
0.01
0.96
Hopper Hop 0.75
0.68
0.01
0.09
Pendulum Swingup 0.75
0.92
0.04
0.03
Walker Run 1.00
Pink
0.01
0.01
Mountain Car 1.25
0.01
0.01
0.45
UMaze Point 2.00
Red
0.01
0.01
0.06
@@
Figure 5 shows the average performance across environments depending on the number of collection environments (x) and noise color (y). A positive association between larger $\beta$ (more correlated noise) and larger number of environments $N_{\textrm{env}}$ becomes visible. This trend is more discernible when the performances for each noise color are compared separately for each number of parallel collection environments. Figure 5 shows the rank of the average performance, separately ranked for each color: a larger number of environments leads to favorable performance of more correlated noise, indicating that with larger samples (i.e., larger $N_{\textrm{env}}$), more exploration is beneficial. However, in combination with the finding from Sec. 3.1, we find the best performing configuration as $\beta=0.5,N_{\textrm{envs}}=4$.
[…]
Figure 5: Preferred noise color depends on number of environments: Average performance across environment, impact of noise color $\beta$ combined with n-envs number of parallel environments: (5) A trend is visible: the averaged performance is larger for larger $\beta$ when more collection environments are used, but the decline due to the number of environments outweighs this trend. (5) Ranks of average performance are indicated by circle size, ranks are calculated across noise-colors but within the same number of environments. The positive trend between number of environments and larger $\beta$ is clearly visible.
@@
(b) Standard deviation of the bias in $\{\varepsilon_{0},\ldots,\varepsilon_{T}\}$, calculated as the standard deviations over the biases in (5(a)) as a function of noise color $\beta$.
@@
We assume that the variance of the bias was optimal for the best performing $\beta$ for each $N_{\textrm{env}}$. We estimate this optimal variance of the bias $\sigma^{*}$ as the average of $\hat{\sigma}(n,\beta)$ for the best performing $\beta$ (Table A2.1). We calculate the difference between all combinations and the optimal combination: ${E_{ij}=(\sigma^{*}-\hat{\sigma}(\beta_{i},N_{j}))^{2}}$ and rank these differences $E_{ij}$ separately for each $N_{\textrm{env}}$ (Figure 7). We find a trend closely matching the performance ranking: larger variance due to larger $\beta$ is compensated by larger number of environments (and thus samples). The theoretical best $\beta$ for each $N_{\textrm{env}}$ (Figure 7) closely follows our empirical observation (Figure 5).
This indicates, that the increased variance in the bias due to larger $\beta$ is compensated by larger $N_{\textrm{env}}$ and since generally more exploration (i.e, larger $\beta$) would be favorable to collect more diverse samples, the set of best trade-off configurations of $(\beta,N_{\textrm{env}})$ shows a positive correlation between $\beta$ and $N_{\textrm{env}}$.
@@ 4 Conclusion
In summary, the results from our large empirical evaluation indicate that $\beta=0.5$ is a better default choice for the noise in PPO. We thus recommend to switch the default noise source choice in PPO to colored noise with $\beta=0.5$.
@@ Acknowledgments
Algorithm 1 Generating Colored Noise using Inverse Fourier Transform
@@
Figure A7.1: Normalized averaged return across environments with $N_{\textrm{env}}=4$ using TRPO: Performance is improved with $\beta=0.5$ also in the TRPO (Schulman et al. 2015) method, providing evidence for the transferability of our results. Due to its popularity the bulk of our investigation focuses on PPO.
@@
Final Return Performance mean max mean max env 4 Rooms Ant 5.50 12.16 3.72 10.50 4 Rooms Point 0.58 1.00 0.40 0.97 4 Rooms Swimmer 22.91 26.22 19.49 26.08 Ball in Cup (Catch) 685.85 966.92 500.22 869.97 Cartpole Balance 742.18 1000.00 608.22 976.19 Cartpole Swingup 126.55 843.10 75.24 627.93 Cheetah Run 222.70 558.55 139.34 351.39 Door 921.96 3128.73 372.29 1937.42 Hopper Hop 5.44 77.86 2.39 34.25 Mountain Car 66.92 98.33 61.88 93.53 Pendulum Swingup 338.21 885.04 205.43 678.91 Reacher Hard 266.71 935.94 139.72 613.31 UMaze Ant 5.49 12.54 3.70 10.78 UMaze Point 0.85 1.00 0.67 0.99 UMaze Swimmer 22.83 26.22 19.47 26.08 Walker Run 105.10 473.11 69.66 194.53
Table A9.1: Return of the final policy (Final Return), mean/max across noise colors and number of parallel collection environments. Return averaged across training (Performance), mean/max across noise colors and number of parallel collection environments
@@
Raffin, Kober, and Stulp (2021) propose to use State Dependent Exploration (SDE). In this experiment we compared SDE exploration in PPO against colored noise. We performed experiments on the same environments as before, using the same number of seeds and variation over $N_{\textrm{env}}$. The default settings for SDE defined in the implementation (Raffin et al. 2021) were used, i.e., SDE was simply turned on by setting use_sde to true. Figure A13.1 shows the result.

### 17 Jun 2024 v2

@@ Abstract
This paper was originally published in: Proceedings of the AAAI Conference on Artificial Intelligence, 38. Available at: https://doi.org/10.1609/aaai.v38i11.29139. This version contains the supplementary material; Code is available at: https://github.com/jkbjh/cn-ppo-paper-code.
@@ 1 Introduction
In this paper, we extend previous work on correlated action-noise in off-policy RL to on-policy RL methods: we introduce and empirically evaluate a modification to PPO incorporating temporally correlated colored noise into the stochastic policy’s distribution. Utilizing the re-parameterization trick, we maintain a Gaussian distributed behavior while injecting correlated noise instead of uncorrelated noise. This introduces a new hyperparameter, the noise color $\beta$ parameterizing the correlated noise.
@@ 1.1 Related Work
In off-policy methods, exploration is often achieved by perturbing the action selection process, for example by perturbing the parameters of the policy (Plappert et al. 2018) or by additively perturbing the action, i.e., by adding action noise. This action perturbation can be done, for example, using temporally uncorrelated Gaussian noise, or temporally correlated noise generated by an Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein 1930). Previous research has shown that increasing the temporal correlation of actions, i.e., switching from Gaussian noise to Ornstein-Uhlenbeck noise, tends to increase state space coverage, but an increase in state space coverage can be beneficial or harmful for policy learning; this depends on the task environment and its dynamics (Hollenstein et al. 2021, 2022). Another method for temporally correlated perturbations was introduced by Rückstieß, Felder, and Schmidhuber (2008); Raffin, Kober, and Stulp (2021) where the “action noise” is generated deterministically by a function only dependent on the state, but the function parameters are randomly changed after a number of steps. This approach is extended by Chiappa et al. (2023) to exploit the hidden layers of the policy, to induce correlation between the action dimensions.
@@ 3.1 Evaluation Details
Environment $\beta^{*}$ Noise Color $\beta\newline 0.5$ $p_{0.5}$ $\beta\newline 0.0$ $p_{0.0}$ $\beta\newline 1.0$ $p_{1.0}$
4 Rooms Swimmer 0.00 White 0.19 0.01
Cartpole Balance 0.00 White 0.04 0.01
Cheetah Run 0.00 White 0.14 0.01
Reacher Hard 0.00 White 0.01 0.01
UMaze Ant 0.00 White 0.24 0.01
UMaze Swimmer 0.00 White 0.16 0.01
4 Rooms Ant 0.20 0.21 0.98 0.01
Door 0.20 0.06 0.56 0.25
Cartpole Swingup 0.50 0.03 0.01
4 Rooms Point 0.75 0.97 0.03 0.51
Ball in Cup (Catch) 0.75 0.90 0.01 0.96
Hopper Hop 0.75 0.68 0.01 0.09
Pendulum Swingup 0.75 0.92 0.04 0.03
Walker Run 1.00 Pink 0.01 0.01
Mountain Car 1.25 0.01 0.01 0.45
UMaze Point 2.00 Red 0.01 0.01 0.06
@@
Figure 5(a) shows the average performance across environments depending on the number of collection environments (x) and noise color (y). A positive association between larger $\beta$ (more correlated noise) and larger number of environments $N_{\textrm{env}}$ becomes visible. This trend is more discernible when the performances for each noise color are compared separately for each number of parallel collection environments. Figure 5(b) shows the rank of the average performance, separately ranked for each color: a larger number of environments leads to favorable performance of more correlated noise, indicating that with larger samples (i.e., larger $N_{\textrm{env}}$), more exploration is beneficial. However, in combination with the finding from Sec. 3.1, we find the best performing configuration as $\beta=0.5,N_{\textrm{envs}}=4$.
(a) Marker size represents averaged performance (scaled to the sixth power to highlight differences)
(b) Marker size represents rank of averaged performance within the same number of environments
Figure 5: Preferred noise color depends on number of environments: Average performance across environment, impact of noise color $\beta$ combined with n-envs number of parallel environments: (5(a)) A trend is visible: the averaged performance is larger for larger $\beta$ when more collection environments are used, but the decline due to the number of environments outweighs this trend. (5(b)) Ranks of average performance are indicated by circle size, ranks are calculated across noise-colors but within the same number of environments. The positive trend between number of environments and larger $\beta$ is clearly visible.
@@
(b) Standard deviation of the bias in $\{\varepsilon_{0},\ldots,\varepsilon_{T}\}$, calculated as the standard deviations over the biases in (6(a)) as a function of noise color $\beta$.
@@
We assume that the variance of the bias was optimal for the best performing $\beta$ for each $N_{\textrm{env}}$. We estimate this optimal variance of the bias $\sigma^{*}$ as the average of $\hat{\sigma}(n,\beta)$ for the best performing $\beta$ (Table A2.1). We calculate the difference between all combinations and the optimal combination: ${E_{ij}=(\sigma^{*}-\hat{\sigma}(\beta_{i},N_{j}))^{2}}$ and rank these differences $E_{ij}$ separately for each $N_{\textrm{env}}$ (Figure 7). We find a trend closely matching the performance ranking: larger variance due to larger $\beta$ is compensated by larger number of environments (and thus samples). The theoretical best $\beta$ for each $N_{\textrm{env}}$ (Figure 7) closely follows our empirical observation (Figure 5(b)).
This indicates, that the increased variance in the bias due to larger $\beta$ is compensated by larger $N_{\textrm{env}}$ and since generally more exploration (i.e., larger $\beta$) would be favorable to collect more diverse samples, the set of best trade-off configurations of $(\beta,N_{\textrm{env}})$ shows a positive correlation between $\beta$ and $N_{\textrm{env}}$.
@@ 4 Conclusion
In summary, the results from our large empirical evaluation indicate that $\beta=0.5$ is a better default choice for the noise in PPO. We thus recommend switching the default noise source choice in PPO to colored noise with $\beta=0.5$.
@@ Acknowledgments
Algorithm 1 Generating Colored Noise using Inverse Fourier Transform.
@@
(a) TRPO
(b) A2C
(c) Overall
Figure A7.1: Normalized averaged return across environments with $N_{\textrm{env}}=4$. Performance is improved with $\beta=0.5$ also in the (a) TRPO method (Schulman et al. 2015), (b) the A2C method (Mnih et al. 2016), providing evidence for the transferability of our results. Due to its popularity the bulk of our investigation focuses on PPO. The effect is less pronounced in (b) A2C, but the overall comparison (c) shows that A2C was not able to learn as well as both PPO and TRPO.
@@
Final Return Performance mean max mean max env 4 Rooms Ant 5.50 12.16 3.72 10.50 4 Rooms Point 0.58 1.00 0.40 0.97 4 Rooms Swimmer 22.91 26.22 19.49 26.08 Ball in Cup (Catch) 685.85 966.92 500.22 869.97 Cartpole Balance 742.18 1000.00 608.22 976.19 Cartpole Swingup 126.55 843.10 75.24 627.93 Cheetah Run 222.70 558.55 139.34 351.39 Door 921.96 3128.73 372.29 1937.42 Hopper Hop 5.44 77.86 2.39 34.25 Mountain Car 66.92 98.33 61.88 93.53 Pendulum Swingup 338.21 885.04 205.43 678.91 Reacher Hard 266.71 935.94 139.72 613.31 UMaze Ant 5.49 12.54 3.70 10.78 UMaze Point 0.85 1.00 0.67 0.99 UMaze Swimmer 22.83 26.22 19.47 26.08 Walker Run 105.10 473.11 69.66 194.53 Table A9.1: Return of the final policy (Final Return), mean/max across noise colors and number of parallel collection environments. Return averaged across training (Performance), mean/max across noise colors and number of parallel collection environments
@@
Raffin, Kober, and Stulp (2021) propose to use generalized State Dependent Exploration (gSDE), an extension over the method proposed by Rückstieß, Felder, and Schmidhuber (2008). In this experiment we compared gSDE exploration in PPO against colored noise. We performed experiments on the same environments as before, using the same number of seeds and variation over $N_{\textrm{env}}$. The default settings for gSDE defined in the implementation by Raffin et al. (2021) were used, i.e., gSDE was simply turned on by setting use_sde to true. Figure A13.1 shows the result.